[Brahms-l] Shift report 20050204 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Feb 04 2005 - 16:01:28 EST
1. At the beginning, it was nice. But around 9:30am, yellow beam has jumped phase. After they dumped the beam around 11:00am, the beam is of quite good quality for data. But they dumped the beam without notification and with Physics ON on BERT around 15:48. 

2. Runs:
12777 - FS@12deg@B1219 - 38k tr.2, MRS@45deg@A2255 - 255k tr.3
12778 - FS@12deg@B861 -  43k tr.6, MRS@45deg@A2255 - 86k  tr.3
(stopped the run for the ZDC rates dropped to 0 suddenly)
12785 - FS@12deg@B861 -  176k tr.6, MRS@45deg@A2255 -  177k tr.3
12786 - FS@12deg@B861 -  148k tr.6, MRS@45deg@A2255 -  146k tr.3
12787 - FS@12deg@B861-ongoing tr.6, MRS@45deg@A2255-ongoing tr.3
Runs 12782(TPC pedestal run)  and 12783  non-TPC pedestal runs.

3. ZDC rates was turned to 0-1 at the beginning(not detector problem in fact), in the new store, everything is back. And Dana fixed ZDC SMX HV problem. HV was not quite connected right on the barrel connector at the PMT.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 4 16:01:48 2005

This archive was generated by hypermail 2.1.8 : Fri Feb 04 2005 - 16:02:06 EST